If i have a formula that requires 16 oz of solution for 75 gallons of water, how much solution do i need for 3 gallons of water
Answer by checkley77(12844)

16/75=X/3
75X=16*3
75X=48
X=48/75
X=.64 OUNCES IS NEEDED.